I am having trouble understanding this question. I don't need every single value worked out I just want to understand what to do so I can work the rest out myself!

The question is written exactly as writtten:

The pK of an acid is 4.16. Calculate the pH of a solution of the acid if it is titrated to the following extents. Use the value associated with your student ID: extent of titration with NaOH, percent

0

1

2

3

4

5

6

7

8

9

28

71

33....

The istructions say,

Show all Calculations, give your pH value to two decimal places

My Student ID number is 1

I really just need help understanding the problem and how I should work it out, I don't want every single pH answer to the solution of the acid if it was titrated to all of the extents that are requested. I truly just want to learn and understand!

Thanks for your help and time!
Answer #1

Given pKa = 4.16

we will use the equation,

pH = pKa + log[salt]/[acid]

when 0% titration is done, all we have is the acid so,

Ka = [H+][A-]/[HA]

let HA = 1 and x be the amount of acid disssociated then,

Ka = x^2/(HA-x), with x being a small amount, the equation reduces to,

Ka = x^2

pKa = -log[Ka]

Ka = 6.92 x 10^-5 = x^2

x = [H+] = 8.32 x 10^-3 M

pH = 2.08

When 5% titration is done,

5% acid has formed salt, 95% is remaining thus,

pH = 4.16 + log(5/95)

      = 2.88

when 50% of titration is done, we would have half of acid reacted to form salt, so,

pH = 4.16 + log(50/50)

pH = 4.16

Similarly other values can be substituted in the equation to calculate the pH.